A piece of sting is wound round a cylinder for 20complete turns.the length of the string is found to be 3.96m.calculate the diameter of the cylinder in cm.
Answer by mananth(16949)


Length of string 3.96 m. = 396 cm
number of rotations = 20
length of string for one rotation = 396/20 = 19.8 cm
pi d =19.2 cm
d=19.6/pi
diameter =6.24 cm